OpenStudy (anonymous):

I need help with solving a problem that I have been trying to figure out for the last 5 hours

OpenStudy (anonymous):

Can I see the problem?

OpenStudy (saifoo.khan):

Lets Try.

OpenStudy (anonymous):

\[\sqrt[3]{1024}\]

OpenStudy (anonymous):

8\[\sqrt[3]{2}\] is not the answer

OpenStudy (anonymous):

\[1024=2^{10}\]

OpenStudy (saifoo.khan):

is it 8 or three?

OpenStudy (anonymous):

according to my answer key it states that the answer is \[4\sqrt[3]{16}\]

OpenStudy (anonymous):

You can still simplify that \[4 \sqrt[3]{2^4}\]

OpenStudy (anonymous):

so: \[4(2)\sqrt[3]{2}\]

OpenStudy (anonymous):

\[8\sqrt[3]{2}\]

OpenStudy (anonymous):

Does that answer you problem?

OpenStudy (anonymous):

not really Josephniel. I have never done cubic sqaures before so i will need to see the entire problem written out and answered with the correct answer.

OpenStudy (anonymous):

cubic roots I meant

OpenStudy (anonymous):

Okay. I'll show you the solution just for the sake of it \[\sqrt[3]{1024}\] Since 1024 is \[ 2^{10} \], we have: \[\sqrt[3]{2^{10}}\] To simplify, since the index is 3, we will divide the exponent to the index. So it's like: \[\ 2^{10/3} \] Doing the operation, we have: \[\ 2^3 * 2^{1/3} \] If your exponent is a fraction, the denominator will serve as the index for your radical sign and the numerator as the exponent for the radicand So, we'll have: \[2^3\sqrt[3]{2}\] \[8\sqrt[3]{2}\]

OpenStudy (anonymous):

The answer in your answer sheet is also right. it's just that the answer is not fully simplified.

OpenStudy (anonymous):

So, satisfied?

OpenStudy (anonymous):

y

OpenStudy (anonymous):

Jose are you on

OpenStudy (anonymous):

yeah.